OpenStudy (anonymous):

if f(x)=cos(logx), then f(x)f(y)-1[f(x)-f(xy)] is

OpenStudy (anonymous):

1[f(x)-f(xy)]... is it 1*[..] or 1/[...]

OpenStudy (anonymous):

Interesting problem: f(xy)=cos(log(xy))=cos(log(x)+log(y))=cos(log(x))cos(log(y))-sin(log(x))sin(log(y)) so f(xy)=f(x)f(y)-sin(log(x))sin(log(y)) now if in the expression you wrote, you ended up subtracting f(xy), the f(x)f(y) would cancel out. But that isn't the way you have it in your expression. Bummer. Maybe it's just as bad as math graffiti can get...or it simplifies algebraically when converting sin(log(x))sin(log(y)) into something in terms of cos(log(x)) and cos(log(y))